# Exploring hidden priors when interpreting gravitational wave and electromagnetic probes of the nuclear equation of state

### v1

@@ II.3 Neutron star mass-radius constraints
NICER observations of periodic X-ray emission on millisecond pulsars provides information about their masses and radii. Though their data product is reported as a weighted posterior distribution, we find these results can be well approximated by a simple Gaussian distribution, as illustrated in Figure 2 for the pulsar PSR-J0030 (in red). In Appendix A we describe the best-fitting gaussian to each NICER (alongside HESS) observation, and the specific method we used to deduce its parameters. In terms of this Gaussian likelihood $L(M,R)$, we can evaluate the marginal likelihood for each equation of state $\alpha$ according to [122, 123, 63]
@@
For comparison, other priors will inherit different priori mass distributions from their prior on the assumed parameterization $s$. If instead one follows Ref. [74] and chooses $s$ to be the central density and $p(\rho_{c})$ to be some prior on that central density, then the the implied mass distribution is proportional to $p(\rho_{c})/dM/d\rho_{c}$. Equivalently, the cumulative distribution function for any individual equation of state satisfies $P(<\rho_{c})=P(<M(\rho_{c}))$, for simplicity assuming a single-branched relationship between mass and central density. The specific prior adopted in [74] very strongly favors massive binaries, where $dR/dM$ is large.
Figure 2: Methodology Mass-Radius likelihoods evaluated for sample cold nuclear EoSs. The pulsar constraints used for the likelihoods are shown in the background in various colors. Darker colors of the M-R curves indicate higher $\log$ likelihood for each EoS. The red and purple patches are the posterior samples from NICER pulse profile of PSR-J0030 and the HESS-J1731 respectively. The horizontal color bands are mass constraints of J0740, J0348, J1614.
Most neutron star mass measurements provide information only about the empirical pulsar star mass distribution, rather than any direct insight into the nuclear equation of state. However, a few very massive pulsars’ existence provides useful insight into the neutron star maximum mass, bounding it from below: J0740 [124, 125], with pulsar mass consistent with a Gaussian with mean and standard deviation $2.14M_{\odot}$ and $0.1M_{\odot}$; J0348+0432 [126], with pulsar mass $2.01M_{\odot}\pm 0.04M_{\odot}$; and J1614 [127], with pulsar mass $1.908M_{\odot}\pm 0.016M_{\odot}$. We follow [28] to express the likelihood of an EoS consistent with these observations as the product of standard normal distributions:
@@ II.4 Gravitational wave observations
Gavitational wave observations of GW170817 constrain the binary masses $m_{i}$ and tidal deformabilities $\Lambda_{i}$ for each component of the binary [1, 2, 3, 128]. Specifically, the presence of matter at leading order impacts the binary’s inspiral, through each components’ dimensionless tidal deformability parameter $\Lambda=(2/3)k_{2}(c^{2}R/GM)^{5}$ [129], where $k_{2}$ is the $l=2$ Love number, $M$ is the mass, and $R$ is the radius. The leading order contribution to the phase evolution of an GW inspiral is given by the weighted combination of $\Lambda$ terms
@@
The impact of tidal effects on outgoing radiation are included in many conventional approximate phenomenological estimates of outgoing radiation from merging compact binaries. In this work, we will focus on analyses with a contemporary state-of-the-art model, IMRPhenomPv2_NRTidalv2 [130, 131], which incorporates precession physics but omits higher-order modes. We note other contemporary models like NRHybSur3dq8Tidal [132] and TEOBResumS [133] include these higher-order modes and tidal effects.
Figure 3 shows our results for the joint posterior in mass, aligned effective inspiral spin $\chi_{\rm eff}$, and tidal deformability $\tilde{\Lambda}$, showing both the posterior (contours) and marginal likelihood (points). For the fiducial analysis (black), the marginal distributions of mass, mass ratio, spin, and $\tilde{\Lambda}$ are consistent with previously published interpretations using comparable models, settings, and data [1, 2]. For comparison and to highlight the impact of astrophysically plausible prior assumptions on source parameter inferences, this figure also shows analyses in which we require either $q>0.9$, $\chi_{i,z}>0$, or both. As the likelihood contains strong correlated features, these constraints impact the posterior distribution of tidal deformability $\tilde{\Lambda}$. When applied individually, these two constraints have modest impact. However, when imposed simultaneously, even these modest constraints produce qualitative changes in the inferred tidal deformability.
Figure 3: GW inference Benchmark inference of properties of GW170817 using RIFT and IMRPhenomP_NRTidalv2. Color scale shows the inferred marginal likelihood, increasing from blue to red over a range of 15 in log-likelihood (with gray indicating values more than $15$ below the peak); 2d contours show the 90% credible intervals; and one-dimensional distributions show the one-dimensional marginal posterior distributions. Parameters shown are the detector-frame chirp mass (denoted here by ${\cal M}_{c}$), mass ratio $q=m_{2}/m_{1}$, aligned effective inspiral spin $\chi_{\rm eff}$, and tidal deformability $\tilde{\Lambda}$. Results for our fiducial analysis (in black) shown here are very similar to published results available through GWOSC [1, 2]. Colors indicate results derived using fiducial priors (black); $q>0.8$ (red); $\chi_{i,z}>0$ (green); and both (blue). The yellow curve shows results derived with a spin prior extending up to $|\chi_{i}|=0.4$.
Figure 4: GW inference assuming high mass ratio A reanalysis of GW170817 using RIFT and IMRPhenomP_NRTidalv2 while requiring $m_{2}/m_{1}<0.6$ As in Figure 3, the color scale shows the marginal likelihood and lines show one or two-dimensional marginal distributions. The solid black line adopts the fiducial spin prior $|\chi_{i}|<0.05$; the red line adopts a prior of similar form but allows much larger component spins $|\chi_{i}|<0.4$. Unlike the previously-described analyses, these investigations use a uniform prior on $\chi_{i,z}$, rather than the “z prior”, to ensure comprehensive coverage. Due to strong correlations between spin and mass ratio, the fiducial spin prior’s arbitrary upper limit strongly constrains the extent of the mass ratio distribution.
Figures 3 and 4 also illustrate inferences derived using even more extreme prior assumptions about the pertinent neutron star spin and mass ratio. For example, for the yellow curve provided in Figure 3, the analysis merely allowed for larger possible values of the NS maximum spin. Relaxing the upper limit on $\chi_{i}$ allows lower-mass $m_{2}$ to remain consistent with the data [2], while increasing $\chi_{i}$ only incrementally. Though this single changed assumption allows for more extreme masses, in isolation this one relaxed assumption doesn’t change our qualitative interpretation; notably, the distribution of $\tilde{\Lambda}$ is nearly unchanged. By contrast, Figure 4 shows inferences derived assuming $m_{2}/m_{1}<0.6$, both adopting fiducial spin magnitudes (black) and larger NS spins (red). Notably, the likelihood includes comparable values for a wide range of mass ratios, including masses $m_{2}<1M_{\odot}$. Contrasting the likelihood (points) to the two posterior marginals (black and green) shows how strongly the arbitrary upper limit in our fiducial spin prior limits the mass ratio distribution. Both extreme-$q$ analyses nonetheless produce comparable $\tilde{\Lambda}$ posterior distributions. Due to correlations between mass ratio and $\tilde{\Lambda}$, the constraint $q<0.6$ causes the posterior distribution of $\tilde{\Lambda}$ to systematically decrease.
@@
To provide a preliminary estimate of the impact of these hidden NS population priors on EoS inference derived from GW170817, we follow Zhao and Lattimer 2018 [145], and adopt simple approximate relationships between tidal deformability and fiducial NS radius $R_{1.4}$. In brief, they argue that to a good approximation $\Lambda\propto(R/M)^{6}$ (Eq. (13) in Ref. [145]); that therefore $\tilde{\Lambda}$ is to a good approximation proportional to $(R_{1.4}/{\cal M}_{c})^{6}$, with a nearly mass-ratio-independent constant (Eq. (15) in Ref. [145]); and therefore for a binary with well-measured chirp mass, the radius of a fiducial NS can be estimated as (evaluated for GW170817)
$R_{1.4}\simeq 13.4{\rm km}(\tilde{\Lambda}/800)^{1/6}\;.$ (10)
Figure 5: GW-derived NS radius Nominal distribution of $R_{1.4}$ derived from Eq. (10) and three GW-only inferences of GW170817 presented in Figures 3, 4: our fiducial analysis (blue); the comparable-mass-ratio positive spin analysis (orange); and the high mass ratio analysis (green).
The three GW-only reanalyses above suggest that hidden astrophysical priors change $\tilde{\Lambda}$ by roughly a factor of 2, increasing the median $\tilde{\Lambda}$ (if we require comparable mass and positive aligned spins) or decreasing it (if we require $q<0.6$) respectively. However, because of the exponent relating $R_{1.4}$ and $\tilde{\Lambda}$, these factors of 2 propagate to a roughly 10% change in the inferred $R_{1.4}$, comparable to the statistical measurement errors usually presented for many of the individual measurements described earlier. Figure 5 shows the cumulative distribution of $R_{1.4}$ derived from this expression and three GW-only analyses: our fiducial analysis (blue); the comparable-mass-ratio positive spin analysis (orange); and the high mass ratio analysis (green).
We emphasize this simple discussion presents radius constraints directly in terms of conventional nonphysical GW priors on $\Lambda_{i}$, treating each as independent and uniformly distributed. These statistical errors are not informed by any EoS model families, nor even by the requirement that both NS have the same EoS. Too, the jointly-uniform prior also produces a marginal prior distribution on $\tilde{\Lambda}$ that goes to zero as $\tilde{\Lambda}$ (i.e., $p(\tilde{\Lambda})\propto\tilde{\Lambda}$ at small $\tilde{\Lambda}$). To facilitate visualization and downstream calculation using posterior samples, other studies have reweighted the GW-only posterior distribution for example to be uniform in $\tilde{\Lambda}$ or $R_{1.4}$; see, e.g., Figure 11 of Ref. [2]. These two priors differ substantially: using Eq. (10), a uniform prior in $\tilde{\Lambda}$ (i.e., $P(<\Lambda)=\Lambda/\Lambda_{max}$) corresponds to a cumulative prior distribution that strongly favors the largest radii (i.e., $P(<R_{1.4})\propto R_{1.4}^{6}$). As our calculations employ a directly-interpolated marginal likelihood $L(X)$, not a smoothed posterior, we do not produce such an intermediate data product. Finally, even though the various model-agnostic GW priors used to construct the posterior distributions shown in Figs. 3, 4 and 5 strongly disfavor extremely small NS radii, the GW likelihood itself does not: very small $\Lambda_{i}$ are nearly as consistent with GW170817 observations as $\Lambda_{i}\simeq O(100)$.
II.5 Combining multiple measurements
@@
We combine multiple real or synthetic observations by multiplying their likelihoods together, along with a prior $p({\cal Y})$ over the nuclear equation of state parameters ${\cal Y}$ (here ${\cal Y}=(\lambda_{o},\lambda_{1}\ldots)$):
@@
In this expression, $n$ indexes the neutron star mass radius observations and $Z_{k}$ is computed according to Eq. (6); $k$ indexes the pulsar mass measurements providing a useful constraint on the maximum neutron star mass, with ${\cal L}_{k}$ computed using one of the factors in Eq. (7); and $g$ indexes the gravitational wave observations, with the GW likelihood computed via Eq. (9). In the above, for simplicity we have not incorporated two factors superfluous to our study but essential to real analyses of the unknown neutron star populations: survey selection biases and models for the unknown neutron star population. For example, unbiased estimates of the nuclear equation of state require self-consistent estimates of the binary neutron star population [106]. In turn, self-consistent population estimates must correctly account for parameter-dependent selection effects effects [146, 147].
@@
To generate the posterior distribution from ${\cal L}_{\rm net}$ and $p({\cal Y})$ we follow the RIFT iterative inference strategy. Starting with some proposed parameters ${\cal Y}_{\alpha}$, we evaluate ${\cal L}_{\rm net}({\cal Y}_{\alpha})\equiv{\cal L}_{net,\alpha}$; build an approximation ^$\hat{\cal L}({\cal Y})$; and use that approximation with an adaptive Monte Carlo integration program to generate both an estimated posterior distribution for ${\cal Y}$ as well as a new set of proposed parameters ${\cal Y}_{\alpha^{\prime}}$ meriting further investigation. Specifically, we have extended the open-source RIFT inference package to include a general-purpose implementation of this algorithm. Our implementation not only provides a physics-neutral implementation of the RIFT algorithm, but also can interface with generic codes which compute marginal evidence factors appearing in multi-event and multi-messenger inference like Eq. (11). Moreover, though not used here as the pertinent factors are constant in our study, our implementation can include the necessary population-averaged integrals needed to account for a flexible population model and survey selection effects associated each class of measurement (i.e., the full inhomogeneous Poisson likelihood [147] or the rate-marginalized likelihood [146, 148]). Similarly, we can also easily account for multinomial leave-few-out averaged results; for example, to marginalize over including the HESS result or not, the factor ${\cal E}$ for the HESS result can be replaced by ${\cal E}p+(1-p)$ and then marginalized over confidence $p$ in that analysis.
@@
Figure 6: Top panel: Likelihood $\ln{\cal Z}$ versus $R_{1.4}$ deduced by comparing realizations of the $\Gamma$ spectral EoS family to the NICER mass-radius constraints of J0030 (blue) and HESS J1731 (orange) Bottom panel: Likelihood $\ln{\cal L}$ versus $M_{max}$. The blue solid, black dashed, and orange dash-dotted lines show Eq. (7), evaluated for each known pulsar mass. The dark-red points show a direct evaluation using realizations of the $\Gamma$ spectral EoS family to the single largest known pulsar mass, for expediency using lalsuite rather than RePrimAnd to evaluate $M_{max}$ for each EoS. The dotted green line shows the hypothesized lower bound on the maximum mass implied by prompt collapse to a black hole; see Section II.7 for discussion.
II.6 Diagnostics
@@
As the posterior distribution is directly proportional to the likelihood ${\cal L}_{\rm net}$, we employ several representations of $\ln{\cal L}_{\rm net}$ and its component factors, to characterize the constraining power of each individual observation. As a concrete example, the top panel in Figure 6 shows our marginal log-likelihood $\ln{\cal Z}$ evaluated using EoS drawn from the full set of marginal likelihood evaluations produced during a hierarchical inference of the EoS, rendered versus $R_{1.4}$, the radius of a fiducial neutron star. In this calculation, the list of equations of state $\alpha$ are drawn from the broad analyses performed below. Even more transparently than the two-dimensional approach adopted in Figure 2, this one-dimensional representation allows us to visually illustrate how well each individual measurement constrains one feature of the equation of state. Similarly, the bottom panel in Figure 6 shows our marginal log-likelihood $\ln{\cal L}$ versus $M_{\rm max}$.
@@
These diagnostics allow us to quickly propagate and interpret the effect of different prior choices about the neutron star population. For example, adopting an alternative neutron star mass distribution that is uniform out to $M_{\rm max}(\alpha)$ shifts log likelihood for each individual factor in $\ln{\cal Z}$ by $-\ln(M_{\rm max}-m_{min})/(2.1M_{\odot}-m_{min})$. Because this systematic factor appears along with every mass-radius measurement, its effects stack as more measurements accumulate, increasingly penalizing configurations with exceptional $M_{\rm max}$ relative to our approach. For expected choices for $M_{\rm max}$, however, this factor is small and should have minimal impact on our posteriors.
II.7 Additional sources of external information
Above, we have highlighted the essential components for any joint analysis for the nuclear EoS from these astrophysical observations, each with relatively well-understood observational and theoretical systematics. However, other groups have proposed employing other sources of information about the nuclear equation of state to further sharpen their analysis. Below, we summarize three particularly tractable approaches to incorporate specific sources of external information.
The nuclear EoS must be consistent with properties of nuclei [97, 73]. Following [45], one strategy for incorporating information about nuclear-density matter is to require the baryon density $n$ and energy density $\rho c^{2}$ to be consistent with constraints on the nuclear symmetry energy $S\simeq\rho/n-m_{n}c^{2}$, assumed to be known to be normally constrained with a mean of $32{\rm MeV}$ and a standard deviation of $2{\rm MeV}$ at saturation density, assumed to be 2.7 ×$\times 10^{14}~{}\rm{g~{}cm^{-3}}$ [74]. Figure 7 shows the marginal likelihood for different EoS realizations, plotted versus two EoS properties $R_{1.4}$ and $M_{\rm max}$. The symmetry energy constraint favors smaller radii at each fixed $M_{max}$.
Based on observations of AT2017gfo, several groups have argued that the observed electromagnetic energetics cannot be consistent with prompt collapse of the remnant to a black hole, using that qualitative constraint to infer an upper bound to the NS maximum mass; see, e.g., [54, 149, 55, 150, 55, 151] and references therein. Roughly speaking,these arguments suggest $M_{max}<2.3M_{\odot}$ [55], extremely close to the mass inferred for J0952-0607 [151].
As noted above, the observation of AT2017gfo favors large amounts of ejecta. While these large ejecta masses can be accommodated with exceptional NS spins or mass ratios, these outcomes can also be produced with large neutron stars. Using this rationale, some groups have argued that the NS tidal deformability should be bounded below by roughly $\tilde{\Lambda}>400$ [53]. Roughly, this proposed constraint on $\tilde{\Lambda}$ has comparable qualitative impact on the information obtained from GW170817 as our constraint that $q>0.9$ and $\chi_{i,z}>0$, as seen in Figure 3. For brevity, we omit further discussion of this proposal.

### v2

@@ II.3 Neutron star mass-radius constraints
NICER observations of periodic X-ray emission on millisecond pulsars provides information about their masses and radii. Though their data product is reported as a weighted posterior distribution, we find these results can be well approximated by a simple Gaussian distribution, as illustrated in Figure 2 for the pulsar PSR-J0030 (red). In Appendix A we describe the best-fitting gaussian to each NICER (alongside HESS) observation, and the specific method we used to deduce its parameters. In terms of this Gaussian likelihood $L(M,R)$, we can evaluate the marginal likelihood for each equation of state $\alpha$ according to [122, 123, 63]
@@
For comparison, other priors will inherit different priori mass distributions from their prior on the assumed parameterization $s$. If instead one follows Ref. [74] and chooses $s$ to be the central density and $p(\rho_{c})$ to be some prior on that central density, then the implied mass distribution is proportional to $p(\rho_{c})/dM/d\rho_{c}$. Equivalently, the cumulative distribution function for any individual equation of state satisfies $P(<\rho_{c})=P(<M(\rho_{c}))$, for simplicity assuming a single-branched relationship between mass and central density. The specific prior adopted in [74] very strongly favors massive binaries, where $dR/dM$ is large.
Figure 2: Methodology Mass-Radius likelihoods evaluated for sample cold nuclear EoSs. The pulsar constraints used for the likelihoods are shown in the background in various colors. Darker colors of the M-R curves indicate higher $\log$ likelihood for each EoS. The EoS with the lowest likelihood is represented with a dashed curve. The red and purple patches are the posterior samples from NICER pulse profile of PSR-J0030 and the HESS-J1731 respectively. The horizontal color bands are mass constraints of J0740, J0348, J1614.
Most neutron star mass measurements provide information only about the empirical pulsar star mass distribution, rather than any direct insight into the nuclear equation of state. However, a few very massive pulsars’ existence provides useful insight into the neutron star maximum mass, bounding it from below: J0740 [124, 125], with pulsar mass consistent with a Gaussian with mean and standard deviation $2.14~{}{\rm M_{\odot}}$ and $0.1~{}{\rm M_{\odot}}$; J0348+0432 [126], with pulsar mass $2.01~{}{\rm M_{\odot}}\pm 0.04~{}{\rm M_{\odot}}$; and J1614 [127], with pulsar mass $1.908~{}{\rm M_{\odot}}\pm 0.016~{}{\rm M_{\odot}}$. We follow [28] to express the likelihood of an EoS consistent with these observations as the product of standard normal distributions:
@@ II.4 Gravitational wave observations
Gravitational wave observations of GW170817 constrain the binary masses $m_{i}$ and tidal deformabilities $\Lambda_{i}$ for each component of the binary [1, 2, 3, 128]. Specifically, the presence of matter at leading order impacts the binary’s inspiral, through each components’ dimensionless tidal deformability parameter $\Lambda=(2/3)k_{2}(c^{2}R/GM)^{5}$ [129], where $k_{2}$ is the $l=2$ Love number, $M$ is the mass, and $R$ is the radius. The leading order contribution to the phase evolution of an GW inspiral is given by the weighted combination of $\Lambda$ terms
@@
The impact of tidal effects are included in many conventional approximate phenomenological estimates of outgoing radiation from merging compact binaries. In this work, we will focus on analyses with a contemporary state-of-the-art model, IMRPhenomPv2_NRTidalv2 [130, 131], which incorporates precession physics but omits higher-order modes. We note other contemporary models like NRHybSur3dq8Tidal [132] and TEOBResumS [133] include these higher-order modes and tidal effects.
@@
To provide a preliminary estimate of the impact of the hidden NS population priors on EoS inference derived from GW170817, we follow Zhao and Lattimer 2018 [136], and adopt simple approximate relationships between tidal deformability and fiducial NS radius $R_{1.4}$. In brief, they argue that to a good approximation $\Lambda\propto(R/M)^{6}$ (Eq. (13) in Ref. [136]); that therefore $\tilde{\Lambda}$ is to a good approximation proportional to $(R_{1.4}/{\cal M}_{c})^{6}$, with a nearly mass-ratio-independent constant (Eq. (15) in Ref. [136]); and therefore for a binary with well-measured chirp mass, the radius of a fiducial NS can be estimated as (evaluated for GW170817)
$R_{1.4}\simeq 13.4~{}{\rm km}(\tilde{\Lambda}/800)^{1/6}\;.$ (10)
II.5 Additional sources of external information
In the preceding subsections, we have highlighted the essential components for any joint analysis for the nuclear EoS from these astrophysical observations, each with relatively well-understood observational and theoretical systematics. However, other groups have proposed employing other sources of information about the nuclear equation of state to further sharpen their analysis. Below, we summarize three particularly tractable approaches to incorporate specific sources of external information.
The nuclear EoS must be consistent with properties of nuclei [98, 73]. Following [45], one strategy for incorporating information about nuclear-density matter is to require the baryon density $n$ and energy density $\rho c^{2}$ to be consistent with constraints on the nuclear symmetry energy $S\simeq\rho/n-m_{n}c^{2}$, assumed to be known to be normally constrained with a mean of $32~{}{\rm MeV}$ and a standard deviation of $2~{}{\rm MeV}$ at saturation density, assumed to be 2.7 ×$\times 10^{14}~{}\rm{g~{}cm^{-3}}$ [74, 137, 138]. Figure 3 shows the marginal likelihood for different EoS realizations, plotted versus two EoS properties $R_{1.4}$ and $M_{\rm max}$. The symmetry energy constraint favors smaller radii at each fixed $M_{max}$.
Based on observations of AT2017gfo, several groups have argued that the observed electromagnetic energetics cannot be consistent with prompt collapse of the remnant to a black hole, using that qualitative constraint to infer an upper bound to the NS maximum mass; see, e.g., [54, 139, 55, 140, 55, 141] and references therein. Roughly speaking, these arguments suggest $M_{max}<2.3~{}{\rm M_{\odot}}$ [55], extremely close to the mass inferred for J0952-0607 [141].
As noted in Section II.2, the observation of AT2017gfo favors large amounts of ejecta. While these large ejecta masses can be accommodated with exceptional NS spins or mass ratios, these outcomes can also be produced with large neutron stars. Using this rationale, some groups have argued that the NS tidal deformability should be bounded below by approximately $\tilde{\Lambda}>400$ [53]. This proposed constraint on $\tilde{\Lambda}$ has comparable qualitative impact on the information obtained from GW170817 as our constraint that $q>0.9$ and $\chi_{i,z}>0$, as will be seen in Section III.2 Figure 7. For brevity, we omit further discussion of this proposal.
A preliminary analysis with the inclusion of a recently measured galactic pulsar J0437 [142] is discussed it in Appendix E.
II.6 Combining multiple measurements
@@
We combine multiple real or synthetic observations by multiplying their likelihoods together, along with a prior $p({\cal Y})$ over the nuclear equation of state parameters ${\cal Y}$ (here ${\cal Y}=(\lambda_{1},\lambda_{2}\ldots)$ the parameters of spectral representation):
@@
In this expression, $n$ indexes the neutron star mass radius observations and $Z_{k}$ is computed according to Eq. (6); $k$ indexes the pulsar mass measurements providing a useful constraint on the maximum neutron star mass, with ${\cal L}_{k}$ computed using one of the factors in Eq. (7); and $g$ indexes the gravitational wave observations, with the GW likelihood computed via Eq. (9). In the equation presented above, for simplicity we have not incorporated two factors superfluous to our study but essential to real analyses of the unknown neutron star populations: survey selection biases and models for the unknown neutron star population. For example, unbiased estimates of the nuclear equation of state require self-consistent estimates of the binary neutron star population [93]. In turn, self-consistent population estimates must correctly account for parameter-dependent selection effects effects [143, 144].
II.7 Hyperparameter posterior pipeline: HyperPipe
To generate the posterior distribution from ${\cal L}_{\rm net}$ and $p({\cal Y})$ we follow the RIFT iterative inference strategy HyperPipe. Starting with some proposed parameters ${\cal Y}_{\alpha}$, we evaluate ${\cal L}_{\rm net}({\cal Y}_{\alpha})\equiv{\cal L}_{net,\alpha}$; build an approximation ^$\hat{\cal L}({\cal Y})$; and use that approximation with an adaptive Monte Carlo integration program to generate both an estimated posterior distribution for ${\cal Y}$ as well as a new set of proposed parameters ${\cal Y}_{\alpha^{\prime}}$ meriting further investigation. This pipeline is able to adaptively explore the space of selected hyperparameters to resolve distinct features in the posterior distribution. The estimate is initiated in a small region of the hyperparameter space (in this case a subsection of the space of $\lambda_{1},\lambda_{2}\ldots$). We refer to the proposed parameter selection, ${\cal Y}_{\alpha}$, as the initial grid of parameters and the resultant physical properties as initial samples. Following the first iteration of evaluating likelihoods a posterior is estimated by performing Monte Carlo integration. Following this new parameters are sampled in the regions of the posterior with lower resolution in order to not exclude important finer details in addition to samples placed in exploratory regions. Following likelihood evaluations for these new points an updated posterior is evaluated by using prior and marginal likelihood information from samples of the previous iteration(s) and the current iteration, hence being better informed. Following this the process is repeated a few times where new sample points are generated and the posterior updated by Monte Carlo integration until the posterior does not deviate between iteration to iteration, resulting in a stable posterior estimate.
@@ II.7 Hyperparameter posterior pipeline: HyperPipe
For this, we have extended the open-source RIFT inference package to include a general-purpose implementation of this algorithm. Our implementation not only provides a physics-neutral implementation of the RIFT algorithm, but also can interface with generic codes which compute marginal evidence factors appearing in multi-event and multi-messenger inference like Eq. (11). Moreover, though not used here as the pertinent factors are constant in this study, our implementation can include the necessary population-averaged integrals needed to account for a flexible population model and survey selection effects associated each class of measurement (i.e., the full inhomogeneous Poisson likelihood [144] or the rate-marginalized likelihood [143, 145]). Similarly, we can also easily account for multinomial leave-few-out averaged results; for example, to marginalize over including the HESS result or not, the factor ${\cal E}$ for the HESS result can be replaced by ${\cal E}p+(1-p)$ and then marginalized over confidence $p$ in that analysis.
@@
Figure 4: Top panel: Likelihood ${\cal Z}$ versus $R_{1.4}$ deduced by comparing realizations of the $\Gamma$ spectral EoS family to the NICER mass-radius constraints of J0030 (blue) and HESS J1731 (orange) Bottom panel: Likelihood ${\cal L}$ versus $M_{max}$. The blue solid, black dashed, and orange dash-dotted lines show Eq. (7), evaluated for each known pulsar mass. The dark-red points show a direct evaluation using realizations of the $\Gamma$ spectral EoS family to the pulsar J0348, for expediency using lalsuite rather than RePrimAnd to evaluate $M_{max}$ for each EoS. The dotted green line shows the hypothesized lower bound on the maximum mass implied by prompt collapse to a black hole; see Section II.5 for discussion.
II.8 Diagnostics
@@
As the posterior distribution is directly proportional to the likelihood ${\cal L}_{\rm net}$, we employ several representations of $\ln{\cal L}_{\rm net}$ and its component factors, to characterize the constraining power of each individual observation. As a demonstration, the top panel in Figure 4 shows our marginal log-likelihood $\ln{\cal Z}$ evaluated using EoS drawn from the full set of marginal likelihood evaluations produced during a hierarchical inference of the EoS, rendered versus $R_{1.4}$, the radius of a fiducial neutron star. In this calculation, the list of equations of state $\alpha$ are drawn from the broad analyses performed in this work. Even more transparently than the two-dimensional approach adopted in Figure 2, this one-dimensional representation allows us to visually illustrate how well each individual measurement constrains one feature of the equation of state. Similarly, the bottom panel in Figure 4 shows our marginal log-likelihood $\ln{\cal L}$ versus $M_{\rm max}$.
@@
These diagnostics allow us to quickly propagate and interpret the effect of different prior choices about the neutron star population. For example, adopting an alternative neutron star mass distribution that is uniform out to $M_{\rm max}(\alpha)$ instead of a fixed value (such as in this study) shifts log likelihood for each individual likelihood contribution in $\ln{\cal Z}$ by a factor $-\ln\frac{M_{\rm max}-m_{min}}{2.1~{}{\rm M_{\odot}}-m_{min}}$. Because this systematic factor appears along with every mass-radius measurement, its effects stack as more measurements accumulate, increasingly penalizing configurations with exceptional $M_{\rm max}$ relative to our approach. For expected choices for $M_{\rm max}$, however, this factor is small and should have minimal impact on our posteriors.